# A topological constraint on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

### v1

@@ 2 Symplectic field theory for exact Lagrangian cobordisms
We fix a $2n$-dimensional simply connected manifold $P$ with a Liouville $1$-form $\lambda$ such that $P$ is the completion of a compact Liouville domain $(\bar{P},\left.\lambda\right|_{\bar{P}})$ satisfying $c_{1}(T\bar{P})=0$. Let us consider
@@
The arguments from Section 2.1 to Section 2.4 are completely parallel to those in [17, Section 3], except the difference of the coefficient rings. In Section 2.5, we will discuss a particular case related to clean Lagrangian intersections.
@@ 2.1 Preliminaries
We fix a complex structure $j_{0}$ on ×$\mathbb{R}\times[0,1]$ such that $(j_{0})_{(s,t)}(\partial_{s})=\partial_{t}$, where $(s,t)$ is the coordinate of ×$\mathbb{R}\times[0,1]$. Let $D\coloneqq\{z\in\mathbb{C}\mid|z|\leq 1\}$. For any $m\in\mathbb{Z}_{\geq 0}$, we fix $p_{0},p_{1},\dots,p_{m}\in\partial D$ ordered counterclockwise and define $D_{m+1}\coloneqq D\setminus\{p_{0},p_{1},\dots,p_{m}\}$. Let $\mathcal{C}_{m+1}$ denote the space of conformal structures on $D_{m+1}$ which can be extended smoothly on $D$. Given $\kappa\in\mathcal{C}_{m+1}$, we let $j_{\kappa}$ denote a complex structure on $D_{m+1}$ which associates $\kappa$. Near each boundary puncture of the Riemann surface $(D_{m+1},j_{\kappa})$, we fix a strip coordinate by a biholomorphic map
@@
Let $\overline{D_{m+1}}$ be a compactification of $D_{m+1}$ obtained by gluing $[0,\infty]\times[0,1]$ via $\psi_{0}$ and $[-\infty,0]\times[0,1]$ via $\psi_{k}$ for every $k=1,\dots,m$. In a natural way, $\overline{D_{m+1}}$ has a structure of a compact oriented surface with boundary and corner.
Let $\Lambda$ be a compact Legendrian submanifold of $(Y,\alpha)$. Later, we will impose the following conditions on $\Lambda$:
  1. 1.
  2. 2.
    For the immersion $\left.\pi_{P}\right|_{\Lambda}\colon\Lambda\to P$, the self-intersection of $\pi_{P}(\Lambda)$ consists only of transversal double points.
  3. 3.
The second condition is satisfied for generic $\Lambda$. We denote the set of Reeb chords of $\Lambda$ by
@@
for every $t\in[0,1]$. When the second condition above is satisfied, the Conley-Zehnder index $\mu(a)\in\mathbb{Z}$ of $a$ is defined by using the capping path $\left.w_{a}\right|_{\partial D_{1}}$ for $a$ as described in [17, Section 3.1]. We then define
$|a|\coloneqq\mu(a)-1.$
@@ 2.2 Moduli space of pseudo-holomorphic curves with boundary in an exact Lagrangian cobordism
Let $\Lambda_{+}$ and $\Lambda_{-}$ be two compact Legendrian submanifolds of $(Y,\alpha)$. Recall that a submanifold $L$ of ×$\mathbb{R}\times Y$ is called an exact Lagrangian cobordism from $\Lambda_{-}$ to $\Lambda_{+}$ if the following are satisfied:
@@
  • There exists a $C^{\infty}$ function $f\colon L\to\mathbb{R}$ such that $df=\left.(e^{r}\alpha)\right|_{L}$ and
@@
    is a constant function. Here, $r$ is the $\mathbb{R}$-coordinate of ×$\mathbb{R}\times Y$.
@@
Let $J$ be an almost complex structure on $P$ such that $d\lambda(\cdot,J\cdot)$ is a Riemannian metric on $P$. We associate to $J$ an almost complex structure $\widetilde{J}$ on ×$\mathbb{R}\times Y$ determined by $\tau_{s}^{*}\widetilde{J}=\widetilde{J}$ for every $s\in\mathbb{R}$, $\widetilde{J}(\operatorname{Ker}\alpha)=\operatorname{Ker}\alpha$$(\pi_{P})_{*}(\widetilde{J}(v))=J((\pi_{P})_{*}(v))$ for every $v\in\operatorname{Ker}\alpha$ and $\widetilde{J}(\partial_{z})=\partial_{r}$.
@@
^$\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$
to be the space of pairs $(u,\kappa)$ of $\kappa\in\mathcal{C}_{m+1}$ and a $C^{\infty}$ map $u\colon D_{m+1}\to P$ such that:
@@
When $m\leq 1$, $\mathcal{C}_{m+1}$ consists of a single element and the group $\mathcal{A}_{m+1}$ of conformal automorphisms on $D_{m+1}$ acts on ^$\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$ by $(u,\kappa)\cdot\varphi\coloneqq(u\circ\varphi,\kappa)$ for every $\varphi\in\mathcal{A}_{m+1}$. As in [17, (10) in Section 3.2], let us define a moduli space to be
@@
In the case where $\Lambda_{+}=\Lambda_{-}=\Lambda$ and $L=\mathbb{R}\times\Lambda$, for any $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$, an $\mathbb{R}$-action on ×$\mathcal{M}_{\mathbb{R}\times\Lambda,(i_{+})_{*}(A)}(a;b_{1},\dots,b_{m})$ is defined by $s\cdot(u,\kappa)\coloneqq(\tau_{s}\circ u,\kappa)$ for every $s\in\mathbb{R}$. Let us denote the quotient space by
¯$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})\coloneqq\mathbb{R}\backslash\mathcal{M}_{\mathbb{R}\times\Lambda,(i_{+})_{*}(A)}(a;b_{1},\dots,b_{m}).$
@@ 2.3 Chekanov-Eliashberg DGA over $\mathbb{Z}[H_{2}(Y,\Lambda)]$
We consider a compact Legendrian submanifold $\Lambda$ of $(Y,\alpha)$ satisfying the three conditions in Section 2.1. We also assume that $\Lambda$ is admissible in the sense of the definition in [6, Section 2.3]. Let us take a spin structure $\mathfrak{s}$ on $\Lambda$.
We take an almost complex structure $J$ on $P$ as in [6, Proposition 2.3] such that the following hold for every $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$: The moduli space ¯$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is cut out transversely and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)-1$. Moreover, when $|b_{1}|+\dots+|b_{m}|=|a|-1$, $\coprod_{A\in H_{2}(Y,\Lambda)}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is a compact 00-dimensional manifold. (For the relation between $J$-holomorphic disks in $P$ defined as in [6, Section 2.3] and $\widetilde{J}$-holomorphic maps in ×$\mathbb{R}\times Y$ defined as in Section 2.2, see [3, Theorem 2.1] and [17, Appendix A.4].)
In addition, ¯$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ has an orientation depending on $\mathfrak{s}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda$. See [17, Definition A.4] for details which follow the conventions in [5] related to orientations.
Let $\mathcal{A}_{*}(\Lambda)$ be the graded $\mathbb{Z}[H_{2}(Y,\Lambda)]$-algebra freely generated by $\mathcal{R}(\Lambda)$ whose graded is given by $|\cdot|\colon\mathcal{R}(\Lambda)\to\mathbb{Z}$. A derivative over $\mathbb{Z}[H_{2}(Y,\Lambda)]$ of degree $-1$
@@
and extended by the Leibniz rule. Here, #$\#_{\operatorname{sign}}$ is the number of the points counted with the signs induced by the orientation.
We refer to [5, Theorem 4.1, Theorem 4.12] which takes $\mathbb{Z}[H_{1}(\Lambda)]$ as the coefficient ring of the Chekanov-Eliashberg DGA. We modify the arguments in their proofs to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda)]$ as the coefficient ring. Then, we have the following theorem.
@@
Let us discuss the effect by changing the spin structure on $\Lambda$. For any two spin structures $\mathfrak{s}_{0}$ and $\mathfrak{s}_{1}$ on $\Lambda$, a cohomology class $d(\mathfrak{s}_{0},\mathfrak{s}_{1})\in H^{1}(\Lambda;\mathbb{Z}/2)$ is defined as in [5, Subsection 4.4.1]. Let $\partial\colon H_{2}(Y,\Lambda)\to H_{1}(\Lambda)$ denote the boundary map. For any $\beta\in H_{1}(\Lambda)$, if we change the choice of a spin structure from $\mathfrak{s}$ to $\mathfrak{s}^{\prime}$, the orientation on the moduli space
@@
is changed by the sign $(-1)^{\langle d(\mathfrak{s},\mathfrak{s}^{\prime}),\beta\rangle}$. See the proof of [5, Theorem 4.29]. We define a ring isomorphism
@@
Then, the next proposition easily follows as in the proof of [5, Theorem 4.30].
@@ 2.4 DGA map associated to an exact Lagrangian cobordism
We consider an exact Lagrangian cobordism $L$ from $\Lambda_{-}$ to $\Lambda_{+}$ as in Section 2.2 such that the Maslov class of $L$ vanishes and both $\Lambda_{+}$ and $\Lambda_{-}$ are connected. We assume that $L$ is oriented and has a spin structure $\mathfrak{s}_{L}$. Then, on a boundary $\Lambda_{+}$ (resp. $\Lambda_{-}$) of $L$, an orientation and a spin structure $\mathfrak{s}_{+}$ (resp. $\mathfrak{s}_{-}$) are induced. See [17, Remark 3.6] for their precise descriptions.
@@
By [8, Lemma 3.8], the moduli space $\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is cut out transversely for generic $J$ and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)$ for any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$. By the Gromov compactness, if $|b_{1}|+\dots+|b_{m}|=|a|$, $\coprod_{C\in H_{2}(\mathbb{R}\times Y,L)}\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is a compact 00-dimensional manifold. In addition, it has an orientation depending on $\mathfrak{s}_{L}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda_{+}$ and $\Lambda_{-}$. See [17, Section A.3].
We refer to the proof of [17, Theorem 3.7] which takes $\mathbb{Z}[H_{1}(\Lambda_{\pm})]$ as the coefficient rings. We modify the arguments in [17, Appendix A.6] to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda_{\pm})]$ as the coefficient rings. (As referred in [17, Appendix A], see also [11, Theorem 2.5] about the DGA map over $\mathbb{Z}$, but note that some conventions in [11] related to orientations are different from those in [5, 17].) Then, we obtain the following theorem.

### v2

@@ 2 Symplectic field theory for exact Lagrangian cobordisms
We fix a $2n$-dimensional simply connected manifold $P$ with a Liouville $1$-form $\lambda$ such that $P$ is the completion of a compact Liouville domain $(\bar{P},\left.\lambda\right|_{\bar{P}})$ satisfying $2c_{1}(T\bar{P})=0$. Let us consider
@@
The arguments from Section 2.1 to Section 2.4 are parallel to those in [16, Section 3], except the difference of the coefficient rings. In Section 2.5, we will discuss a particular case related to clean Lagrangian intersections.
@@ 2.1 Preliminaries
We fix a complex structure $j_{0}$ on $\mathbb{R}\times[0,1]$ such that $(j_{0})_{(s,t)}(\partial_{s})=\partial_{t}$, where $(s,t)$ is the coordinate of $\mathbb{R}\times[0,1]$. Let $D\coloneqq\{z\in\mathbb{C}\mid|z|\leq 1\}$. For any $m\in\mathbb{Z}_{\geq 0}$, we fix $p_{0},p_{1},\dots,p_{m}\in\partial D$ ordered counterclockwise and define $D_{m+1}\coloneqq D\setminus\{p_{0},p_{1},\dots,p_{m}\}$. Let $\mathcal{C}_{m+1}$ denote the space of conformal structures on $D_{m+1}$ which can be extended smoothly on $D$. Given $\kappa\in\mathcal{C}_{m+1}$, we let $j_{\kappa}$ denote a complex structure on $D_{m+1}$ which associates $\kappa$. Near each boundary puncture of the Riemann surface $(D_{m+1},j_{\kappa})$, we fix a strip coordinate by a biholomorphic map
@@
Let $\overline{D_{m+1}}$ be a compactification of $D_{m+1}$ obtained by gluing $0,\infty]\times[0,1]$ via $\psi_{0}$ and $\infty,0]\times[0,1]$ via $\psi_{k}$ for every $k=1,\dots,m$. In a natural way, $\overline{D_{m+1}}$ has a structure of a compact oriented surface with boundary and corners.
Let $\Lambda$ be a compact Legendrian submanifold of $(Y,\alpha)$. We denote the set of Reeb chords of $\Lambda$ by
@@
for every $t\in[0,1]$.
@@ 2.2 Moduli space of pseudo-holomorphic curves with boundary in an exact Lagrangian cobordism
Let $\Lambda_{+}$ and $\Lambda_{-}$ be two compact Legendrian submanifolds of $(Y,\alpha)$. Recall that a submanifold $L$ of $\mathbb{R}\times Y$ is called an exact Lagrangian cobordism from $\Lambda_{-}$ to $\Lambda_{+}$ if the following hold:
@@
  • There exists a $C^{\infty}$ function $f\colon L\to\mathbb{R}$ such that $df=\left.(e^{r}\alpha)\right|_{L}$ and its restriction
@@
    is a constant function. Here, $r$ is the $\mathbb{R}$-coordinate of $\mathbb{R}\times Y$.
@@
Let $J$ be an almost complex structure on $P$ such that $d\lambda(\cdot,J\cdot)$ is a Riemannian metric on $P$ and adapted to $\Lambda$ in the sense of the definition in [5, Section 2.3]. We associate to $J$ an almost complex structure $\widetilde{J}$ on $\mathbb{R}\times Y$ determined by $\tau_{s}^{*}\widetilde{J}=\widetilde{J}$ for every $s\in\mathbb{R}$, $\widetilde{J}(\operatorname{Ker}\alpha)=\operatorname{Ker}\alpha$, $(\pi_{P})_{*}(\widetilde{J}(v))=J((\pi_{P})_{*}(v))$ for every $v\in\operatorname{Ker}\alpha$ and $\widetilde{J}(\partial_{z})=\partial_{r}$.
@@
$\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$
to be the space of pairs $(u,\kappa)$ of $\kappa\in\mathcal{C}_{m+1}$ and a $C^{\infty}$ map $u\colon D_{m+1}\to\mathbb{R}\times Y$ such that:
@@
When $m\leq 1$, $\mathcal{C}_{m+1}$ consists of a single element and the group $\mathcal{A}_{m+1}$ of conformal automorphisms on $D_{m+1}$ acts on $\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$ by $(u,\kappa)\cdot\varphi\coloneqq(u\circ\varphi,\kappa)$ for every $\varphi\in\mathcal{A}_{m+1}$. Let us define a moduli space to be
@@
In the case where $\Lambda_{+}=\Lambda_{-}=\Lambda$ and $L=\mathbb{R}\times\Lambda$, for any $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$, an $\mathbb{R}$-action on $\mathcal{M}_{\mathbb{R}\times\Lambda,(i_{+})_{*}(A)}(a;b_{1},\dots,b_{m})$ is defined by $s\cdot(u,\kappa)\coloneqq(\tau_{s}\circ u,\kappa)$ for every $s\in\mathbb{R}$. Let us denote the quotient space by
$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})\coloneqq\mathbb{R}\backslash\mathcal{M}_{\mathbb{R}\times\Lambda,(i_{+})_{*}(A)}(a;b_{1},\dots,b_{m}).$
@@ 2.3 Chekanov-Eliashberg DGA over $\mathbb{Z}[H_{2}(Y,\Lambda)]$
We consider a compact Legendrian submanifold $\Lambda$ of $(Y,\alpha)$ satisfying the following conditions:
  1. 1.
  2. 2.
    For the immersion $\left.\pi_{P}\right|_{\Lambda}\colon\Lambda\to P$, the self-intersection of $\pi_{P}(\Lambda)$ consists only of transverse double points.
  3. 3.
The second condition is satisfied for generic $\Lambda$. Under the second condition, the Conley-Zehnder index $\mu(a)\in\mathbb{Z}$ of $a$ is defined by using the capping path $\left.w_{a}\right|_{\partial D_{1}}$ for $a$ as in [16, Section 3.1]. We then define
$|a|\coloneqq\mu(a)-1.$
We also assume that $\Lambda$ is admissible in the sense of the definition in [5, Section 2.3]. Let us fix a spin structure $\mathfrak{s}$ on $\Lambda$.
We take an almost complex structure $J$ on $P$ as in [5, Proposition 2.3] such that the following hold for every $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$ such that $(b_{1},\dots,b_{m})\neq(a)$: The moduli space $\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is cut out transversely and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)-1$. Moreover, when $|b_{1}|+\dots+|b_{m}|=|a|-1$, $\coprod_{A\in H_{2}(Y,\Lambda)}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is a compact 0-dimensional manifold. (For the relation between $J$-holomorphic disks in $P$ defined as in [5, Section 2.3] and $\widetilde{J}-$holomorphic maps in $\mathbb{R}\times Y$ defined as in Section 2.2, see [2, Theorem 2.1] and [16, Appendix A.4].)
In addition, $\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ has an orientation depending on $\mathfrak{s}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda$. See [16, Definition A.4] for details which follow the conventions in [4] related to orientations.
Let $\mathcal{A}_{*}(\Lambda)$ be the graded $\mathbb{Z}[H_{2}(Y,\Lambda)]$-algebra freely generated by $\mathcal{R}(\Lambda)$ whose grading is given by $|\cdot|\colon\mathcal{R}(\Lambda)\to\mathbb{Z}$. A derivation over $\mathbb{Z}[H_{2}(Y,\Lambda)]$ of degree 1-1
@@
and extended by the Leibniz rule. Here, $_{\operatorname{sign}}$ is the number of the points counted with the signs induced by the orientation.
We refer to [4, Theorem 4.1, Theorem 4.12] which takes $\mathbb{Z}[H_{1}(\Lambda)]$ as the coefficient ring of the Chekanov-Eliashberg DGA. We modify the arguments in their proof to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda)]$ as the coefficient ring. Then, we obtain the following theorem.
@@
Let us discuss the effect by changing the spin structure on $\Lambda$. For any two spin structures $\mathfrak{s}_{0}$ and $\mathfrak{s}_{1}$ on $\Lambda$, a cohomology class $d(\mathfrak{s}_{0},\mathfrak{s}_{1})\in H^{1}(\Lambda;\mathbb{Z}/2)$ is defined as in [4, Subsection 4.4.1]. Let $\partial\colon H_{2}(Y,\Lambda)\to H_{1}(\Lambda)$ denote the boundary map. For any $\beta\in H_{1}(\Lambda)$, if we change the choice of a spin structure from $\mathfrak{s}$ to $\mathfrak{s}^{\prime}$, the orientation on the moduli space
@@
is changed by the sign $(-1)^{\langle d(\mathfrak{s},\mathfrak{s}^{\prime}),\beta\rangle}$. See the proof of [4, Theorem 4.29]. We define a ring isomorphism
@@
Then, the next proposition holds as in the proof of [4, Theorem 4.30].
@@ 2.4 DGA map associated to an exact Lagrangian cobordism
We consider an exact Lagrangian cobordism $L$ from $\Lambda_{-}$ to $\Lambda_{+}$ as in Section 2.2 such that the Maslov class of $L$ vanishes and both $\Lambda_{+}$ and $\Lambda_{-}$ are connected. We assume that $L$ is oriented and has a spin structure $\mathfrak{s}_{L}$. Then, on positive end $\Lambda_{+}$ (resp. the negative end $\Lambda_{-}$) of $L$, an orientation and a spin structure $\mathfrak{s}_{+}$ (resp. $\mathfrak{s}_{-}$) are induced. See [16, Remark 3.6] for their precise descriptions.
@@
By [7, Lemma 3.8], the moduli space $\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is cut out transversely for generic $J$ and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)$ for any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$. By the Gromov compactness, if $|b_{1}|+\dots+|b_{m}|=|a|$, $\coprod_{C\in H_{2}(\mathbb{R}\times Y,L)}\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is a compact 0-dimensional manifold. In addition, it has an orientation depending on $\mathfrak{s}_{L}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda_{+}$ and $\Lambda_{-}$. See [16, Section A.3].
We refer to the proof of [16, Theorem 3.7] which takes $\mathbb{Z}[H_{1}(\Lambda_{\pm})]$ as the coefficient rings. We modify the arguments in [16, Appendix A.6] to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda_{\pm})]$ as the coefficient rings. (As referred in [16, Appendix A], see also [10, Theorem 2.5] about the DGA map over $\mathbb{Z}$, but note that some conventions in [10] related to orientations are different from those in [4, 16].) Then, we obtain the following theorem.